OpenStudy (anonymous):

Please I'm failing math and need to understand this?Will fan and medal please! You earn $17.50 an hour and work 40 hours a week. your deductions are FICA 7.65%, federal tax withholdings 12.3%, state tax withholding 6.2%. your housing and fixed expenses are 30% of your realized income per month. you want to save 5 months worth in an emergency fund within a year. how much do you need to save each month to fund the emergency fund, and how much discretionary money remains per month? Thanks!!

OpenStudy (anonymous):

@TINY10207

OpenStudy (anonymous):

im so sorry i will try to find out the answer

OpenStudy (anonymous):

I know 17.50/hr*40/hr=$700/wk but i dont know how to take out the deductions

OpenStudy (anonymous):

@dan815 @Famous5

OpenStudy (anonymous):

the total income = 17.5*40=700 weekly = 700*4=2800 monthly total deductions % = 7.65+12.3+ 6.2 + 30= 56.15% emergency fund = 700*4*5= 14,000 annually %= 14000/12=1166.67 monthly =2800*56.15/100=1572.2 emergency fund % =1166.67/2800 *100 = 41.66 %

OpenStudy (anonymous):

so you need to save 41.66 per month and the remaining discretionary money is 1,572.20?

OpenStudy (anonymous):

@AJ01

OpenStudy (anonymous):

No. discretionary income amount that you stated is annually, am i right?

OpenStudy (anonymous):

Your Q says Monthly.

OpenStudy (anonymous):

Plus, thats emergancy fund not discretionary

OpenStudy (anonymous):

no its asking for discretionary monthly. i know the emergency isnt discretionary but its asking how much you need to save each month and how much discretionary is left each month.

OpenStudy (anonymous):

This is the discretionary income monthly: $405.53

OpenStudy (anonymous):

You make $1572.2 monthly after all deductions you've stated, correct? Subtract that with $1166.67 emergency fund you need to save each month. This equates that amount you can spend on other things.

OpenStudy (anonymous):

What do you think?

OpenStudy (anonymous):

You need to save $1166.67 each month?

OpenStudy (anonymous):

for emergency fund ya.

OpenStudy (anonymous):

or if you prefer 'fund' instead of save

OpenStudy (anonymous):

I say save b/c its your invest for the future

OpenStudy (anonymous):

okay that just seems a bit high 700*4*5= 14,000 why is the 5 there?

OpenStudy (anonymous):

Ill figure it out for you in a bit. brb. P.S. i did not write that.

OpenStudy (anonymous):

Back. Explanation: $700 (= weekly salary) x 4 (= 4 weeks in a month (normally)) x 5 (= 5 months worth of emergency fund)

OpenStudy (anonymous):

This equates the amount of emergency fund for 5 months based on your monthly income $2,800.

OpenStudy (anonymous):

So each month you would save?

OpenStudy (anonymous):

@7k7

OpenStudy (anonymous):

Remember: you want to save 5 months worth in an emergency fund within a year. Save: Emergency Fund per year = $14,000, So per month would be $1166.67 for a year.

OpenStudy (anonymous):

'Save' per month: 1. Emergency Fund $1,166.67 2. discretionary money $405.53

OpenStudy (anonymous):

Okay thank you!

OpenStudy (anonymous):

Because... $1,572.2 - $1,166.67 = $405.53. Subtract the monthly emergency fund ($1,166.67) from your monthly income after all deductions ($1,572.2) to see how much you have left to spend.

OpenStudy (anonymous):

Your Welcome!